OpenStudy (anonymous):

log__1/16=-2 I'm trying to find the base

OpenStudy (anonymous):

suppose base is b b^-2 = 1/16

OpenStudy (mertsj):

So basically you are trying to solve this equation: \[x ^{-2}=\frac{1}{16}\]

OpenStudy (anonymous):

b^-2 = 4^-2 therefore b = 4

OpenStudy (anonymous):

Oh okay I how you all solved this.. Thank you so much

OpenStudy (anonymous):

Okay how will you go about log_6=-1/3